# Reasoning Runtime Behavior of a Program with LLM: How Far Are We?

### v2

@@ VI-A Case Study
Figure 5 shows a case from the problem of EPP. Given a Python function that aims to return the largest prime factor of parameter n and input $15$, this problem requires the model to predict the next statement to be executed after an assignment statement (i.e., largest = max ...). Here we select three models that show competitive performance in EPP including GPT-4-Turbo, GPT-3.5-Turbo, and CodeLlama-34B-Instruct. For GPT-3.5-Turbo and CodeLlama-34B-Instruct, the prediction is the last “return” statement of this function; while GPT-4-Turbo chooses the above “for” loop as its prediction. We marked a happy emoji to present that GPT-4-Turbo makes the right choice, and the other two models fail to predict it correctly. The explanation for the answer is that if the assignment statement is executed, the possible value of $j$ can only be three or five (n + 1 = 16), which means the loop will go on and the “return” statement is not the next executed line.
In our framework and the adapted benchmark, there are many problems like this that have no obvious answer and are challenging. If the model is not capable of reasoning about its inherent logic of execution, it can be easily misled and give the most “look-alike” answer (i.e., the “return” statement in this case), which reflects that our framework can effectively measure the code reasoning capability of LLMs and present the discrimination of them.
@@ VI-B Unsatisfactory Performance of Code Reasoning
VI-C Threats to Validity
Internal Threats. To construct the adapted benchmark for different sub-tasks of Runtime Behavior Reasoning, we manually establish some rules to select appropriate statements and variables for the evaluation. However, our selection criteria may not effectively represent the runtime state of the program. To mitigate this threat, we take some measures to determine problem settings that are representative and challenging, based on the characteristics of different tasks. For instance, we choose the last statement in the control flow (i.e., for EPP) and variables that are modified after the execution (i.e., for PSP). These measures help us to reasonably assess the model’s capability to reason about code and provide meaningful differentiation. For Runtime Behavior Reasoning, we select four dimensions of the intermediate state of program execution which are widely applied in previous research [16, 38]. These four tasks are proven to effectively evaluate the code reasoning capability of code LLMs [38, 16, 15], and they are appropriate for Incremental Consistency Evaluation for their unique sequential relationship. However, there are still some dynamic features such as memory allocation and exception handling which may help measure code models and we have not explored yet. Further research could consider exploring the potential for LLMs to reason about other dynamic program features and extend $\bm{\mathcal{R}}$Eval to more scenarios.
External Threats. In the empirical study for our evaluation framework, the results are restricted to the specific collection of code models and base benchmarks. To mitigate this threat, we choose the representative code LLMs considering several standards including their scale, popularity, and training strategy; For base benchmarks applied, two benchmarks (i.e., HumanEval [8] and ClassEval [22]) are distinct from evaluation fashion and programming scenarios. With the above efforts, the experimental results are expected to be held under more circumstances.
@@ VII Conclusion and Future Work
In this paper, we propose $\bm{\mathcal{R}}$Eval, a comprehensive framework for evaluating the code reasoning capability of code LLMs. Our framework consists of two evaluation components including Runtime Behavior Reasoning and Incremental Consistency Evaluation: We conduct a large-scale empirical study on several popular LLMs and two widely used base benchmarks. Our empirical results show that the majority of LLMs we evaluate show unsatisfactory performance on both Runtime Behavior Reasoning and Incremental Consistency Evaluation. To improve the code reasoning capabilities of LLMs, future works can explore:
Training with Execution Behavior. One reason why large models may struggle with code reasoning is possibly due to a lack of knowledge related to program execution. Although some general fine-tuning approaches are applied to code LLMs [7], they fail to improve the code reasoning capabilities. Given the demonstrated effectiveness of training models with execution behavior in enhancing performance across a range of downstream tasks [16, 23], it is reasonable to expect that LLMs would also derive benefits from such a process.

### v3

@@ VI-A Case Study
Fig. 5 shows a case from the problem of EPP. Given a Python function that aims to return the largest prime factor of parameter n and input $15$, this problem requires the model to predict the next statement to be executed after an assignment statement (i.e., largest = max ...). Here we select three models that show competitive performance in EPP including GPT-4-Turbo, GPT-3.5-Turbo, and CodeLlama-34B-Instruct. For GPT-3.5-Turbo and CodeLlama-34B-Instruct, the prediction is the last “return” statement of this function; while GPT-4-Turbo chooses the above “for” loop as its prediction. We mark a happy emoji to show that GPT-4-Turbo makes the right choice, and the other two models fail to predict it correctly. The explanation is that if the assignment statement is executed, the possible value of $j$ can only be 3 or 5 ($n+1=16$), which means that the loop will continue and the “return” statement is not the next executed line.
In our framework and the adapted benchmark, there are many problems like this that have no obvious answer and are challenging. If the model is not capable of reasoning about its inherent logic of execution, it can easily be misled and give the most “look-alike” answer (i.e., the “return” statement in this case), indicating that our framework can effectively measure the code reasoning capability of LLMs and present the discrimination of them. We discuss more cases in the appendix, which can be accessed in our replication package [^1] https://r-eval.github.io.
@@ VI-B Unsatisfactory Performance of Code Reasoning
TABLE IV: Pearson Correlation Coefficient Matrix of the Results of Runtime Behavior Reasoning (RBR), Incremental Consistency (IC), and HumanEval (HE)
Pearson Correlation RBR IC HE
RBR 1.000 0.940 0.772
IC 0.940 1.000 0.724
HE 0.772 0.724 1.000
VI-C Correlation between Code Reasoning & Code Generation
We utilize the experimental results and study the correlation between code reasoning and code generation, i.e., whether an LLM that performs well in code generation could exhibit equally strong abilities in code reasoning.
Table IV presents the Pearson correlation coefficient matrix of Runtime Behavior Reasoning (Avg. Acc.), Incremental Consistency Score and HumanEval (pass@$1$ rate). According to the matrix, we find that there is a strong positive correlation (i.e, the Pearson correlation coefficients are larger than 0.7) among code generation (i.e., HumanEval) and code reasoning (i.e., Runtime Behavior Reasoning and Incremental Consistency). However, the correlation between code reasoning and code generation is relatively lower than that between two reasoning tasks internally (i.e., 0.724 and 0.772 v.s. 0.940), which indicates that models with similar code generation abilities may vary a lot in code reasoning. As the correlation may help researchers increase the understanding of code LLMs and improve the models’ code generation and reasoning abilities, future research could investigate such correlation in depth.
VI-D Threats to Validity
Internal Threats. To construct the adapted benchmark for different sub-tasks of Runtime Behavior Reasoning, we manually establish some rules to select appropriate statements and variables for the evaluation. However, our selection criteria may not effectively represent the runtime state of the program. To mitigate this threat, we take some measures to determine problem settings that are representative and challenging, based on the characteristics of different tasks. For instance, we choose the last statement in the control flow (i.e., for EPP) and variables that are modified after the execution (i.e., for PSP). These measures help us to reasonably assess the model’s capability to reason about code and provide meaningful differentiation. For Runtime Behavior Reasoning, we select four dimensions of the intermediate state of program execution which are widely applied in previous research [16, 42]. These four tasks are proven to effectively evaluate the code reasoning capability of code LLMs [42, 16, 15], and are appropriate for Incremental Consistency Evaluation for their unique sequential relationship. However, there are still some dynamic features such as memory allocation and exception handling which may help measure code models, and we have not explored yet. Further research could consider exploring the potential for LLMs to reason about other dynamic program features and extend $\bm{\mathcal{R}}$Eval to more scenarios.
External Threats. In the empirical study for our evaluation framework, the results are restricted to the specific collection of code models and base benchmarks. To mitigate this threat, we choose representative code LLMs considering several standards including their scale, popularity, and training strategy; For the base benchmarks applied, two benchmarks (i.e., HumanEval [8] and ClassEval [22]) are distinct from evaluation fashion and programming scenarios, as described in Section III-D. With the above efforts, the experimental results are expected to be sustained in more circumstances.
@@ VII Conclusion and Future Work
In this paper, we propose $\bm{\mathcal{R}}$Eval, a comprehensive framework for evaluating the code reasoning capability of code LLMs. Our framework consists of two evaluation components including Runtime Behavior Reasoning and Incremental Consistency Evaluation: We conduct a large-scale empirical study on several popular LLMs and two widely used base benchmarks. Our empirical results show that the majority of LLMs we evaluate show unsatisfactory performance in both Runtime Behavior Reasoning and Incremental Consistency Evaluation. To improve the code reasoning capabilities of LLMs, future works can explore:
Training with Execution Behavior. One reason why large models may struggle with code reasoning is possibly due to a lack of knowledge related to program execution. Although some general fine-tuning approaches are applied to code LLMs [7], they fail to improve the code reasoning capabilities. Given the demonstrated effectiveness of training models with execution behavior in improving performance in a range of downstream tasks [16, 24], it is reasonable to expect that LLMs would also derive benefits from such a process.
@@
Acknowledgments
This research is supported by the Ningbo Natural Science Foundation (No. 2023J292). It is also supported by the advanced computing resources provided by the Supercomputing Center of Hangzhou City University.